OpenStudy (anonymous):

What would log8 7 – log8 s + log8 t – log8 4 be simplified? please explain

OpenStudy (anonymous):

let y = log8 7 - log8 s + log8 t - log8 4 8^y = 8^(log8 7 - log8 s + log8 t - log8 4) 8^y = 8^(log8 7) * 8^(-log8 s) * 8^(log8 t) * 8^(-log8 4) 8^y = 7*(1/s)*(t)*(1/4) = 7t/4s log8 8^y = y = log8 (7t/4s)

OpenStudy (anonymous):

how would this one be done? Simplify 7 log3 k + 6 log3 m – 9 log3 n